# LLM Meets Scene Graph: Can Large Language Models Understand and Generate Scene Graphs? A Benchmark and Empirical Study

### v1

@@ Abstract
Dongil Yang1    Minjin Kim1    Sunghwan Kim1    Beong-woo Kwak1 Minjun Park1    Jinseok Hong2    Woontack Woo2    Jinyoung Yeo1thanks: Corresponding author Yonsei University1 KAIST2 {wingu,jinyeo}@yonsei.ac.kr
@@ 2 Related Work
Scene graphs, which encode semantic information and relationships between objects in a scene, are widely used in various tasks across vision, language, and multimodal domains such as image captioning, 3D reconstruction, games and robotics. (Johnson et al., 2015; Yao et al., 2018; Chatterjee et al., 2021; Armeni et al., 2019; Ammanabrolu and Riedl, 2021; Ammanabrolu and Hausknecht, 2020; Rosinol et al., 2020). Previous research has leveraged scene graphs as an inductive bias to facilitate explicit reasoning in visual question answering tasks (Teney et al., 2017; Hildebrandt et al., 2020) or to provide explanations based on such knowledge (Shi et al., 2018). Scene graphs can also serve as semantic maps of real-world environments for robots, enabling higher-level reasoning for tasks like planning and navigation (Rana et al., 2023; Dai et al., 2023; Yin et al., 2024).
@@ 3.4 Dataset Construction
We additionally construct data for our understanding tasks through a similar collaboration process. For SGQA, we provide the LLM with the entire scenario narrative and prompt it to generate five questions about identifying a node that had undergone spatial and temporal transitions. Human workers then verify the validity of these questions. For SGDS, we control the difficulty of the distractors by perturbing the answer description in two ways. We put random distractors from unrelated scenarios for the half, and put distractors with LLM-perturbed nodes and edges from the answer for the other half of the problems. More detailed descriptions of our dataset construction process are in Appendix A.
@@ 5.4 Can LLMs Refine Errors in Scene Graphs?
5.5 Hallucination in Scene Graph Generation
Hallucination is a known issue in LLMs. We investigate hallucination in the scene graph generation task. If the LLM generates elements not included in the predefined elements $L$, they are counted as hallucinations.
@@ 5.5 Hallucination in Scene Graph Generation
Qwen-2.5-7B 395 192 203
Table 6: Hallucination counts for each model, showing the total number, those semantically similar to the words in the description, and those entirely unrelated. “Desc.” indicates “Description Elements.” and “New.” indicates “New Elements.”
Although hallucinations occur, their absolute number is low. We categorize them into two types: (1) those caused by generating elements semantically similar to elements mentioned in the description, and (2) those arising from producing entirely new elements. Table 1 shows that larger models mostly produce the former, while smaller models tend to generate both, indicating higher susceptibility.
We further test the impact of decoding temperature. We set the temperature to 1.0, and under otherwise identical conditions, we observed a slight increase in hallucinations for both nodes and edges.
In addition to observing an increased rate of hallucinations in both nodes and edges, we also evaluate the effect on performance in multi-action scene graph generation. GPT-4o’s performance drops from 43.99 to 39.72, and Sonnet’s from 58.8 to 56.8. These results suggest that higher temperature can increase hallucinations and degrade structured output quality.
@@ Appendix A Dataset Construction Details
Then, we generate scene graphs with GPT-4o based on each narrative resulted from Step 1. Here, we let the model generate elements from the predefined set of the EASG dataset. As the model often fails to generate complete graphs, human workers check each instance one by one, whether any mismatch between the description and scene graph exists, including missing elements. During this process, we add some frequently appearing elements and obtain the set of valid nodes and edges for each scenario.
@@
Although we have annotations for description and scene graph pairs, we additionally work on improving the quality of the text narratives, in terms of lexical diversity, coherency, and complexity. We prompt GPT-4o to paraphrase each sentence, given the preceding context, and also identify which word changes it has made. As this process may produce a description containing a similar element in the predefined set, which may yield a misleading annotation with the reference graph, we skip the instance. After paraphrasing, we prompt the model again to increase fluency regarding the context (e.g., adding a conjunction or modifier).
Data construction for understanding tasks.
For generating questions for SGQA, we provide GPT-4o with the entire context in graph form to make multi-hop questions. We prompt it to make questions about common elements across multiple triplets (e.g., first object that had been grabbed), the state of nodes under specific conditions (e.g., before or after an action happened), etc.. For SGDS, we also prompt GPT-4o to perturb the candidates to generate sentences with similar words but that have different meanings with the reference. To finalize, human reviewers inspect and correct these questions and answers to ensure accuracy.
@@
To ensure consistency and quality across annotations, we implemented a structured alignment phase at the early stage of data construction. All annotators participated in a sampling-based alignment session under the supervision of a project lead. During this process, annotators were required to follow a set of detailed checklists specific to each step of the annotation workflow. These checklists served as shared standards to minimize variance and ensure high-quality outputs.
@@
  • Only values from the predefined scenario-specific list (denoted as $L$) may be used for nodes and edges.
@@
  • All major actions and relations described must be reflected in the scene graph without omission, and LLM-generated graphs must be thoroughly reviewed and supplemented by annotators.
@@
  • The meaning of the original sentence must be preserved in paraphrased versions.
  • Synonym replacements must not result in multiple correct answers for evaluation purposes
(e.g., replacing place with put-down should not introduce ambiguity with other valid verb forms like lay or set.)
Paraphrased sentences must still be representable with the original scene graph format.
When merging multiple sentences, the resulting sentence must comprehensively cover all the corresponding scene graphs.
If the merged sentence includes information that can be logically inferred, explicit mention may be omitted.
@@ Appendix C Additional Analysis
C.1 Navigating Edge Formation Under Varied Information Levels
We examined how sentence length and edge density influence generation quality. In lengthy sentences with few relations, LLMs can often filter out irrelevant details while maintaining precision. However, short sentences with numerous implied relationships (e.g., hand usages, repeated actions) provide limited explicit cues, requiring the model to infer missing links. While LLMs manage noisy sentences reasonably well, their performance weakens when extensive inference is needed with minimal textual guidance.
Figure 7: The results show $F_{1}$ performance between multiple action scene graph($\mathit{M}$) and single action scene graph($\mathit{S}$) approaches across description-to-edge ratio ranges. The histogram represents sample frequencies in each ratio range.
C.2 Case Study for Descriptions with Repetitive Action
The failure cases are shown in Table  C.2.

### v2

@@ Abstract
Dongil Yang1    Minjin Kim1    Sunghwan Kim1    Beong-woo Kwak1 Minjun Park1    Jinseok Hong2    Woontack Woo2, 3    Jinyoung Yeo1thanks: Corresponding author Department of Artificial Intelligence, Yonsei University 1 Graduate School of Metaverse, KAIST2 Graduate School of Culture Technology, KAIST3 {wingu,jinyeo}@yonsei.ac.kr
@@ 2 Related Work
Scene graphs, which encode semantic information and relationships between objects in a scene, are widely used in various tasks across vision, language, and multimodal domains such as image captioning, 3D reconstruction, interactive games and robotics. (Johnson et al., 2015; Yao et al., 2018; Chatterjee et al., 2021; Armeni et al., 2019; Ammanabrolu and Riedl, 2021; Ammanabrolu and Hausknecht, 2020; Rosinol et al., 2020). Previous research has leveraged scene graphs as an inductive bias to facilitate explicit reasoning in visual question answering tasks (Teney et al., 2017; Hildebrandt et al., 2020) or to provide explanations based on such knowledge (Shi et al., 2018). Scene graphs can also serve as semantic maps of real-world environments for robots, enabling higher-level reasoning for tasks like planning and navigation (Rana et al., 2023; Dai et al., 2023; Yin et al., 2024).
@@ 3.4 Dataset Construction
We additionally construct data for our understanding tasks through a similar collaboration process. For SGQA, we provide the LLM with the entire scenario narrative and prompt it to generate five questions about identifying a node that had undergone spatial and temporal transitions. Human workers then verify the validity of these questions. For SGDS, we control the difficulty of the distractors by perturbing the answer description in two ways. We put random distractors from unrelated scenarios for the half, and put distractors with LLM-perturbed nodes and edges from the answer for the other half of the problems. More detailed descriptions of our dataset construction process are in Appendix 6.
@@ 5.4 Can LLMs Refine Errors in Scene Graphs?
5.5 Do LLMs Hallucinate in Scene Graph Generation?
Since hallucination is a known issue of LLMs, we investigate whether they hallucinate in the SA-SGG task. We consider hallucination occurred when any generated element is not present in the predefined set of elements, $L$. To account for instances in which the model simply copies elements that are explicitly mentioned in the provided description $d_{i}$, we further classify hallucinated outputs into two categories: (1) elements that are mentioned in the description but not included in $L$, and (2) entirely new elements not present in either $L$ or the description. As shown in Table 6, Claude-3.5-Sonnet exhibits a very low incidence of hallucination, whereas smaller models tend to hallucinate more frequently.
@@ 5.5 Do LLMs Hallucinate in Scene Graph Generation?
Table 6: Hallucination counts in the SA-SGG task(1,188 data). "Total" indicates the total number of hallucinations, “Desc.” indicates “Description Elements,” which are explicit in the description but not in $L$, regarded as related but incorrect elements, and “New.” indicates “New Elements,” which are not included in either $L$ or the description, regarded as entirely unrelated elements.
@@ Appendix A Dataset Construction Details
Then, we generate scene graphs with GPT-4o based on each narrative resulted from Step 1. Here, we let the model generate the graph, guided by the predefined elements of the EASG dataset. The generated elements are then grouped by scenario (with each scenario corresponding to multiple scene graphs). Then, we let human workers review each instance, consisting of a text description, a scene graph, and scenario-specific elements. Following this review process, the scene graph and the scenario-specific elements—denoted as $L$—are finalized as the reference data.
@@
Although we have annotations for description and scene graph pairs, we additionally work on improving the quality of the text narratives, in terms of lexical diversity, coherency, and complexity. We prompt GPT-4o to paraphrase each sentence, given the preceding context, and also identify which word changes it has made. As this process may produce a description containing a similar element in the scenario-specific elements, which may yield a misleading annotation with the reference graph, we skip the instance. After paraphrasing, we prompt the model again to increase fluency regarding the context (e.g., adding a conjunction or modifier).
@@
To ensure consistency and quality across annotators, we underwent an alignment phase before each step. All annotators participated in a sampling-based alignment session under the supervision of a project lead. During this process, they were required to follow a set of detailed checklists specific to each step of the annotation workflow. These checklists served as shared standards to minimize variance and ensure high-quality outputs.
@@
  • All elements must be lemmatized.
@@
  • All major actions and relations described must be reflected in the scene graph without omission.
@@
  • If synonyms are present in the predefined set of scenario-specific elements, only one representative term should be retained.
  • The set of scenario-specific elements may be expanded with additional elements in the graph, if they are deemed necessary based on the context.
@@
  • If the meaning of the original sentence is not preserved in paraphrased versions, the original sentence should be retained.
  • If the paraphrasing produces any different element that is present in the scenario-specific set, the original sentence should be retained, in order to avoid ambiguous annotation.
  • If the paraphrased sentence is not representable with the target scene graph format, the original sentence should be retained.
  • When multiple sentences are merged, the resulting sentence must comprehensively cover all the corresponding scene graphs. Unless, the original sentence should be retained.
  • Even if multiple sentences are implicitly expressed as a single sentence, it’s acceptable as long as the original sentences can be logically inferred.
Data construction for understanding tasks.
For generating questions for SGQA, we provide GPT-4o with the entire context in graph form to make multi-hop questions. We prompt it to make questions about common elements across multiple triplets (e.g., first object that had been grabbed), the state of nodes under specific conditions (e.g., before or after an action happened), etc.. For SGDS, we also prompt GPT-4o to perturb the candidates to generate sentences with similar words but that have different meanings with the reference. To finalize, human reviewers manually edit any incorrect question or answer to ensure accuracy.
Figure 7: The line graph shows $F_{1}$ scores of GPT-40 and Claude across varying levels of descriptiveness in the MA-SGG data. The histogram represents sample frequencies within each descriptiveness ratio range.
@@ Appendix C Additional Analysis
C.1 Analysis of the Impact of Description Length on Scene Graph Generation
We examined the effect of the descriptiveness of $d_{i}$ on scene graph generation performance. Using the MA-SGG data, we quantified descriptiveness as the ratio of the number of words in each description $d_{i}$ to the number of edges in the corresponding scene graphs $G_{i}$. A lower ratio indicates that the description is written concisely, whereas a higher ratio reflects a description with greater detail to express the scene. As shown in Figure 7, both GPT-4o and Claude-3.5-Sonnet perform better when provided with more detailed descriptions, which likely offer additional cues for constructing accurate scene graphs.
C.2 Case Study for the MA-SGG Task
As LLMs show low performance in the MA-SGG task, we show some failure cases in Table  C.2.